A 2.00-g sample of SX (g) has a volume of 329.5 cm at 1.00 atm and 20 °C. Identify element X and name the compound.
Answer #1

Volume (V) = 329.5 cm3 = 0.3295 Liters

Pressure (P) = 1 atm

Temperature (T) = 20 oC = 293.15 K

Gas constant (R) = 0.0821 L atm mol-1 K-1

PV = nRT

No. of moles (n) = PV / RT = (1 atm x 0.3295 L) / (0.0821 L atm mol-1 K-1 x 293.15 K) = 0.0137 mol

Moles = weight / molar mass

molar mass = weight / moles = (2 g) / (0.0137 mol) = 146 g/mol

Given compound is: SX6

(32 g/mol) + (6 x molar mass of X) = 146 g/mol

molar mass of X = (114 / 6) = 19 g/mol

molar mass having 19 g/mol = Fluorine

So;

the compound is: SF6